\begin{thebibliography}{126}

\bibitem{MUSES:2023hyz}
R.~Kumar et~al. (MUSES), Living Rev. Rel. \textbf{27}, 3 (2024), \texttt{2303.17021}.

\bibitem{Stephanov:1998dy}
M.A. Stephanov, K.~Rajagopal, E.V. Shuryak, Phys. Rev. Lett. \textbf{81}, 4816 (1998).

\bibitem{Bollweg:2022rps}
D.~Bollweg, {{et al.}} (HotQCD), Phys. Rev. D \textbf{105}, 074511 (2022).

\bibitem{Fu:2019hdw}
W.j. Fu, {{et al.}}, Phys. Rev. D \textbf{101}, 054032 (2020).

\bibitem{Gunkel:2021oya}
P.J. Gunkel, C.S. Fischer, Phys. Rev. D \textbf{104}, 054022 (2021), \texttt{2106.08356}.

\bibitem{Basar:2021hdf}
G.~Basar, Phys. Rev. Lett. \textbf{127}, 171603 (2021), \texttt{2105.08080}.

\bibitem{Dimopoulos:2021vrk}
P.~Dimopoulos, {{et al.}}, Phys. Rev. D \textbf{105}, 034513 (2022).

\bibitem{Hippert:2023bel}
M.~Hippert, {{et al}} (2023), \texttt{2309.00579}.

\bibitem{Sorensen:2024mry}
A.~Sorensen, P.~Sorensen (2024), \texttt{2405.10278}.

\bibitem{ZHANGSQM}
Y.~ZHANG, This {SQM2024} proceedings (2024)

\bibitem{Bzdak:2019pkr}
A.~Bzdak, {{et al}}, Phys. Rept. \textbf{853}, 1 (2020).

\bibitem{Dore:2022qyz}
T.~Dore, {{et al}}, Phys. Rev. D \textbf{106}, 094024 (2022).

\bibitem{Nonaka:2004pg}
C.~Nonaka, M.~Asakawa, Phys. Rev. C \textbf{71}, 044904 (2005), \texttt{nucl-th/0410078}.

\bibitem{Pradeep:2024cca}
M.S. Pradeep, N.~Sogabe, M.~Stephanov, H.U. Yee, Phys. Rev. C \textbf{109}, 064905 (2024).

\bibitem{Karthein:2024zvs}
J.M. Karthein, V.~Koch, C.~Ratti (2024), \texttt{2409.13961}.

\bibitem{Borsanyi:2021sxv}
S.~Bors\'anyi, {{et al}}, Phys. Rev. Lett. \textbf{126}, 232001 (2021).

\bibitem{Kahangirwe:2024cny}
M.~Kahangirwe, {{et al}}, Phys. Rev. D \textbf{109}, 094046 (2024).

\bibitem{Plumberg:2024leb}
C.~Plumberg et~al. (2024), \texttt{2405.09648}.

\bibitem{MUSESalpha}
{{MUSES}} calculation enginge (2024), \urlstyle{tt}\url{https://musesframework.io/}

\bibitem{Shen:2017bsr}
C.~Shen, B.~Schenke, Phys. Rev. C \textbf{97}, 024907 (2018), \texttt{1710.00881}.

\bibitem{Cimerman:2023hjw}
J.~Cimerman, {{et al}}, Phys. Rev. C \textbf{107}, 044902 (2023).

\bibitem{Garcia-Montero:2023gex}
O.~Garcia-Montero, {{et al}}, Phys. Rev. C \textbf{109}, 044916 (2024).

\bibitem{Garcia-Montero:2024jev}
O.~Garcia-Montero, S.~Schlichting (2024), \texttt{2409.06788}.

\bibitem{Schafer:2021csj}
A.~Sch\"afer, {{et al}} (SMASH), Eur. Phys. J. A \textbf{58}, 230 (2022).

\bibitem{Carzon:2019qja}
P.~Carzon, {{et al}}, Phys. Rev. C \textbf{105}, 034908 (2022).

\bibitem{Carzon:2023zfp}
P.~Carzon, {{et al}}, Phys. Rev. C \textbf{108}, 064905 (2023).

\bibitem{DoreSQM}
T.~Dore, This {SQM2024} proceedings (2024)

\bibitem{Monnai:2024pvy}
A.~Monnai, G.~Pihan, B.~Schenke, C.~Shen (2024), \texttt{2406.11610}.

\bibitem{Karpenko:2013wva}
I.~Karpenko, {{et al}}, Comput. Phys. Commun. \textbf{185}, 3016 (2014).

\bibitem{Almaalol:2022pjc}
D.~Almaalol, T.~Dore, J.~Noronha-Hostler (2022), \texttt{2209.11210}.

\bibitem{Kapusta:2024riv}
J.~Kapusta, M.~Singh, T.~Welle (2024), \texttt{2409.16525}

\bibitem{Karthein:2024ipl}
J.M. Karthein, {{et al}} (2024), \texttt{2409.16249}

\bibitem{An:2022jgc}
X.~An, G.~Basar, M.~Stephanov, H.U. Yee, Phys. Rev. C \textbf{108}, 034910 (2023).

\bibitem{Pradeep:2022eil}
M.S. Pradeep, M.~Stephanov, Phys. Rev. Lett. \textbf{130}, 162301 (2023), \texttt{2211.09142}.

\bibitem{Inghirami:2022afu}
G.~Inghirami, H.~Elfner, Eur. Phys. J. C \textbf{82}, 796 (2022), \texttt{2201.05934}.

\bibitem{Sorensen:2020ygf}
A.~Sorensen, V.~Koch, Phys. Rev. C \textbf{104}, 034904 (2021), \texttt{2011.06635}.

\bibitem{Bedaque:2014sqa}
P.~Bedaque, A.W. Steiner, Phys. Rev. Lett. \textbf{114}, 031103 (2015), \texttt{1408.5116}.

\bibitem{Tan:2020ics}
H.~Tan, {{et al}}, Phys. Rev. Lett. \textbf{125}, 261104 (2020), \texttt{2006.16296}.

\bibitem{Kojo:2020krb}
T.~Kojo, AAPPS Bull. \textbf{31}, 11 (2021), \texttt{2011.10940}.

\bibitem{McLerran:2018hbz}
L.~McLerran, S.~Reddy, Phys. Rev. Lett. \textbf{122}, 122701 (2019), \texttt{1811.12503}.

\bibitem{Cruz-Camacho:2024odu}
N.~Cruz-Camacho, {{et al}} (2024).

\bibitem{Tan:2021ahl}
H.~Tan, {{et al}}, Phys. Rev. D \textbf{105}, 023018 (2022).

\bibitem{Essick:2023fso}
R.~Essick, {{et al}}, Phys. Rev. D \textbf{108}, 043013 (2023).

\bibitem{Mroczek:2023zxo}
D.~Mroczek, M.C. Miller, J.~Noronha-Hostler, N.~Yunes (2023), \texttt{2309.02345}.

\bibitem{Alford:2013aca}
M.G. Alford, S.~Han, M.~Prakash, Phys. Rev. D \textbf{88}, 083013 (2013), \texttt{1302.4732}.

\bibitem{Tan:2021nat}
H.~Tan, {{et al}}, Phys. Rev. Lett. \textbf{128}, 161101 (2022).

\bibitem{Brandt:2022hwy}
B.B. Brandt, F.~Cuteri, G.~Endrodi, JHEP \textbf{07}, 055 (2023), \texttt{2212.14016}.

\bibitem{Komoltsev:2021jzg}
O.~Komoltsev, A.~Kurkela, Phys. Rev. Lett. \textbf{128}, 202701 (2022), \texttt{2111.05350}.

\bibitem{Most:2021zvc}
E.R. Most, {{et al}}, Mon. Not. Roy. Astron. Soc. \textbf{509}, 1096 (2021).

\bibitem{Alford:2020pld}
M.G. Alford, A.~Haber, Phys. Rev. C \textbf{103}, 045810 (2021), \texttt{2009.05181}.

\bibitem{Yang:2023ogo}
Y.~Yang, M.~Hippert, E.~Speranza, J.~Noronha, Phys. Rev. C \textbf{109}, 015805 (2024).

\bibitem{Ripley:2023lsq}
J.L. Ripley, A.~Hegade K.~R., R.S. Chandramouli, N.~Yunes (2023), \texttt{2312.11659}.

\bibitem{Oliinychenko:2022uvy}
D.~Oliinychenko, {{et al}}, Phys. Rev. C \textbf{108}, 034908 (2023).

\bibitem{Yao:2023yda}
N.~Yao, {{et al}}, Phys. Rev. C \textbf{109}, 065803 (2024).

\bibitem{Mroczek:2024sfp}
D.~Mroczek, {{et al}} (2024), \texttt{2404.01658}.

\bibitem{Raithel:2019gws}
C.A. Raithel, F.~Ozel, D.~Psaltis, Astrophys. J. \textbf{875}, 12 (2019), \texttt{1902.10735}.

\bibitem{Heinz:2013th}
U.~Heinz, R.~Snellings, Ann. Rev. Nucl. Part. Sci. \textbf{63}, 123 (2013), \texttt{1301.2826}.

\bibitem{Bernhard:2019bmu}
J.E. Bernhard, J.S. Moreland, S.A. Bass, Nature Phys. \textbf{15}, 1113 (2019).

\bibitem{JETSCAPE:2020mzn}
D.~Everett et~al. (JETSCAPE), Phys. Rev. C \textbf{103}, 054904 (2021), \texttt{2011.01430}.

\bibitem{Nijs:2020roc}
G.~Nijs, {{et al}}, Phys. Rev. C \textbf{103}, 054909 (2021).

\bibitem{Plumberg:2021bme}
C.~Plumberg, {{et al}}, Phys. Rev. C \textbf{105}, L061901 (2022).

\bibitem{Domingues:2024pom}
T.S. Domingues, {{et al}} (2024).

\bibitem{PHENIX:2018lia}
C.~Aidala et~al. (PHENIX), Nature Phys. \textbf{15}, 214 (2019), \texttt{1805.02973}.

\bibitem{STAR:2017ckg}
L.~Adamczyk et~al. (STAR), Nature \textbf{548}, 62 (2017), \texttt{1701.06657}.

\bibitem{Florkowski:2017ruc}
W.~Florkowski, B.~Friman, A.~Jaiswal, E.~Speranza, Phys. Rev. C \textbf{97}, 041901 (2018).

\bibitem{Gursoy:2014aka}
U.~Gursoy, D.~Kharzeev, K.~Rajagopal, Phys. Rev. C \textbf{89}, 054905 (2014), \texttt{1401.3805}.

\bibitem{Niemi:2015voa}
H.~Niemi, K.J. Eskola, R.~Paatelainen, K.~Tuominen, Phys. Rev. C \textbf{93}, 014912 (2016).

\bibitem{Noronha-Hostler:2015uye}
J.~Noronha-Hostler, M.~Luzum, J.Y. Ollitrault, Phys. Rev. C \textbf{93}, 034912 (2016).

\bibitem{ALICE:2016ccg}
J.~Adam et~al. (ALICE), Phys. Rev. Lett. \textbf{116}, 132302 (2016), \texttt{1602.01119}.

\bibitem{Bally:2022vgo}
B.~Bally et~al. (2022), \texttt{2209.11042}.

\bibitem{Giacalone:2023cet}
G.~Giacalone, G.~Nijs, W.~van~der Schee, Phys. Rev. Lett. \textbf{131}, 202302 (2023).

\bibitem{Hu:2021trw}
B.~Hu et~al., Nature Phys. \textbf{18}, 1196 (2022), \texttt{2112.01125}.

\bibitem{PREX:2021umo}
D.~Adhikari et~al. (PREX), Phys. Rev. Lett. \textbf{126}, 172502 (2021), \texttt{2102.10767}.

\bibitem{Summerfield:2021oex}
N.~Summerfield, {{et al}}, Phys. Rev. C \textbf{104}, L041901 (2021).

\bibitem{Giacalone:2024luz}
G.~Giacalone et~al. (2024), \texttt{2402.05995}.

\bibitem{Giacalone:2024ixe}
G.~Giacalone et~al. (2024), \texttt{2405.20210}.

\bibitem{ALICE491854}
ALICE, Ali-prel-491854

\bibitem{Zhao:2022ayk}
W.~Zhao, C.~Shen, B.~Schenke, Phys. Rev. Lett. \textbf{129}, 252302 (2022), \texttt{2203.06094}.

\bibitem{Floerchinger:2021ygn}
S.~Floerchinger, {{et al}}, Phys. Rev. C \textbf{105}, 044908 (2022).

\bibitem{Inghirami:2019mkc}
G.~Inghirami, {{et al}}, Eur. Phys. J. C \textbf{80}, 293 (2020).

\bibitem{Inghirami:2016iru}
G.~Inghirami, {{et al}}, Eur. Phys. J. C \textbf{76}, 659 (2016).

\bibitem{Mayer:2024kkv}
M.~Mayer, L.~Rezzolla, H.~Elfner, G.~Inghirami, D.H. Rischke (2024), \texttt{2403.08669}.

\bibitem{Most:2021rhr}
E.R. Most, J.~Noronha, Phys. Rev. D \textbf{104}, 103028 (2021), \texttt{2109.02796}.

\bibitem{Bemfica:2020xym}
F.S. Bemfica, {{et al}}, Phys. Rev. Lett. \textbf{126}, 222301 (2021).

\bibitem{Cordeiro:2023ljz}
I.~Cordeiro, {{et al}}, Phys. Rev. Lett. \textbf{133}, 091401 (2024).

\bibitem{Abboud:2023hos}
N.~Abboud, E.~Speranza, J.~Noronha, Phys. Rev. D \textbf{109}, 094007 (2024), \texttt{2308.02928}.

\bibitem{Daher:2024bah}
A.~Daher, {{et al}}, Phys. Rev. D \textbf{109}, 114001 (2024).

\bibitem{PalermoSQM}
A.~Palermo, This {SQM2024} proceedings (2024)

\bibitem{Andronic:2017pug}
A.~Andronic, {{et al}}, Nature \textbf{561}, 321 (2018).

\bibitem{Bellwied:2018tkc}
R.~Bellwied, {{et al}}, Phys. Rev. C \textbf{99}, 034912 (2019).

\bibitem{Alba:2020jir}
P.~Alba, {{et al}}, Phys. Rev. C \textbf{101}, 054905 (2020).

\bibitem{Borsanyi:2014ewa}
S.~Borsanyi, {{et al}}, Phys. Rev. Lett. \textbf{113}, 052301 (2014).

\bibitem{Floris:2014pta}
M.~Floris, Nucl. Phys. A \textbf{931}, 103 (2014), \texttt{1408.6403}.

\bibitem{Bazavov:2014xya}
A.~Bazavov et~al., Phys. Rev. Lett. \textbf{113}, 072001 (2014), \texttt{1404.6511}.

\bibitem{SanMartin:2023zhv}
J.S. San~Martin, {{et al}} (2023), \texttt{2309.01737}.

\bibitem{Bellwied:2013cta}
R.~Bellwied, {{et al}}, Phys. Rev. Lett. \textbf{111}, 202302 (2013).

\bibitem{Andronic:2018qqt}
A.~Andronic, {{et al}}, Phys. Lett. B \textbf{792}, 304 (2019).

\bibitem{Vovchenko:2024wbg}
V.~Vovchenko, Phys. Rev. C \textbf{110}, 034914 (2024), \texttt{2405.16306}.

\bibitem{ZherebtsovaSQM}
E.~Zherebtsova, This {SQM2024} proceedings (2024)

\bibitem{Brylinski:2023nrb}
W.~Brylinski, {{et al}} (2023).

\bibitem{Kapusta:2023xrw}
J.I. Kapusta, S.~Pratt, M.~Singh, Phys. Rev. C \textbf{109}, L031902 (2024), \texttt{2306.13280}.

\bibitem{BassSQM}
S.~Bass, This {SQM2024} proceedings (2024)

\bibitem{Capellino:2022nvf}
F.~Capellino, {{et al}}, Phys. Rev. D \textbf{106}, 034021 (2022).

\bibitem{Capellino:2023cxe}
F.~Capellino, {{et al}}, Phys. Rev. D \textbf{108}, 116011 (2023).

\bibitem{HirayamaSQM}
R.~Hirayama, This {SQM2024} proceedings (2024)

\bibitem{Cao:2018ews}
S.~Cao et~al., Phys. Rev. C \textbf{99}, 054907 (2019), \texttt{1809.07894}.

\bibitem{Nahrgang:2014vza}
M.~Nahrgang, {{et al}}, Phys. Rev. C \textbf{91}, 014904 (2015).

\bibitem{Noronha-Hostler:2016eow}
J.~Noronha-Hostler, {{et al}}, Phys. Rev. Lett. \textbf{116}, 252301 (2016).

\bibitem{Prado:2016szr}
C.A.G. Prado, {{et al}}, Phys. Rev. C \textbf{96}, 064903 (2017).

\bibitem{Sambataro:2022sns}
M.L. Sambataro, {{et al}}, Eur. Phys. J. C \textbf{82}, 833 (2022).

\bibitem{Barreto:2022ulg}
L.~Barreto, {{et al}} (2022), \texttt{2208.02061}.

\bibitem{Holtermann:2023vwr}
A.~Holtermann, {{et al}}, Phys. Rev. C \textbf{108}, 064901 (2023).

\bibitem{Sambataro:2023tlv}
M.L. Sambataro, V.~Minissale, S.~Plumari, V.~Greco, Phys. Lett. B \textbf{849}, 138480 (2024).

\bibitem{Cho:2023kpe}
S.~Cho, Phys. Rev. C \textbf{109}, 054904 (2024), \texttt{2307.14765}.

\bibitem{ATLAS:2021ktw}
G.~Aad et~al. (ATLAS), Phys. Rev. C \textbf{105}, 064903 (2022), \texttt{2111.06606}.

\bibitem{Zhao:2024ecc}
J.~Zhao, {{et al}}, Phys. Rev. C \textbf{110}, 024909 (2024).

\bibitem{Karmakar:2024jak}
B.~Karmakar, {{et al}} (2024), \texttt{2403.17817}.

\bibitem{CMS:2018loe}
A.M. Sirunyan et~al. (CMS), Phys. Rev. Lett. \textbf{121}, 082301 (2018), \texttt{1804.09767}.

\bibitem{CMS:2018duw}
A.M. Sirunyan et~al. (CMS), Phys. Lett. B \textbf{791}, 172 (2019), \texttt{1810.01473}.

\bibitem{CMS:2023dse}
A.~Tumasyan et~al. (CMS), Phys. Lett. B \textbf{850}, 138518 (2024), \texttt{2310.03233}.

\bibitem{ATLAS:2023vms}
G.~Aad et~al. (ATLAS), Phys. Rev. Lett. \textbf{132}, 202301 (2024), \texttt{2308.16652}.

\bibitem{Katz:2019fkc}
R.~Katz, {{et al}}, Phys. Rev. C \textbf{102}, 024906 (2020).

\bibitem{Minissale:2023dct}
V.~Minissale, {{et al}}, Eur. Phys. J. C \textbf{84}, 228 (2024).

\bibitem{Faraday:2023mmx}
C.~Faraday, A.~Grindrod, W.A. Horowitz, Eur. Phys. J. C \textbf{83}, 1060 (2023).

\bibitem{Moore:2004tg}
G.D. Moore, D.~Teaney, Phys. Rev. C \textbf{71}, 064904 (2005), \texttt{hep-ph/0412346}.

\bibitem{Pooja:2024rnn}
Pooja, M.Y. Jamal, P.P. Bhaduri, M.~Ruggieri, S.K. Das (2024), \texttt{2404.05315}.

\end{thebibliography}

\documentclass{webofc}

\usepackage[varg]{txfonts}   
\usepackage{hyperref}
\usepackage{url}
\hypersetup{colorlinks=true,citecolor=blue,urlcolor=blue,linkcolor=blue}
\begin{document}
\title{State-of-the-art of Strangeness in Quark Matter Theory 2024}
%
%

\author{\firstname{Jacquelyn} \lastname{Noronha-Hostler}\inst{1}\fnsep\thanks{\email{jnorhos@illinois.edu}}
}

\institute{Illinois Center for Advanced Studies of the Universe, Department of Physics, 
University of Illinois at Urbana-Champaign, 1110 W. Green St., Urbana IL 61801-3080, USA
}

\abstract{
I discuss the theoretical developments in  related to Strangeness in Quark Matter (SQM) leading up to the SQM2024 conference. These advances include mapping out the Quantum Chromodynamics phase diagram; puzzles that exist in hadron physics from light to heavy particles; and advanced in relativistic hydrodynamics with the inclusion of spin and magnetic fields.  
}
\maketitle
\section{Introduction}
\label{intro}
 I will summarize recent advances in high-energy nuclear theory, pressing questions within the field,  and make connections to astrophysics, gravity, nuclear structure,  and cold atoms.

\section{QCD Phase Diagram}
\label{sec:QCDphases}

The QCD phase diagram cannot be systematically studies in a single experiment or astronomy observation, rather it requires input from heavy-ion collisions HIC (high temperatures $T$, low to intermediate $n_B$, range of charge fractions $Y_Q=Z/A=n_Q/n_B=[0.38,0.5]$), cold neutron stars NS ($T=0$, large $n_B$, $Y_Q\lesssim 0.1$), binary NS mergers ($T>0$, large $n_B$, $Y_Q\sim  [0.01,0.2]$), and low energy nuclear physics ($T\sim 0$, nuclear saturation density $n_B=n_{sat}$, $Y_Q\sim 0.5$). 
These leave gaps in the QCD phase diagram that may be difficult to interpret.  
Theoretical input is important e.g. causality and stability constraints $0\leq c_s^2\leq 1$,  first principle lattice or perturbative QCD (pQCD) calculations, or well-defined effective field theories (EFT) such as chiral EFT. 
Constraints are summarized in \cite{MUSES:2023hyz}.
\emph{Key questions: } How to interpret  new Beam Energy Scan II data? How does QCD matter move at $n_B\neq 0$?   How does HIC/QCD constrain the NS equation of state (EOS)?  Can out-of-equilibrium effects shed light on the core of NS?

{\bf HIC/critical point} The (potential) QCD critical point \cite{Stephanov:1998dy}  separates the boundary between quarks/gluons and hadrons. 
While current first principle calculations cannot determine its existence/location \cite{Bollweg:2022rps}, recent calculations in other theoretical frameworks \cite{Fu:2019hdw,Gunkel:2021oya,Basar:2021hdf,Dimopoulos:2021vrk,Hippert:2023bel,Sorensen:2024mry}  indicate a critical point at $\mu_B/T\sim 4-8$. 
Preliminary measurements exist for net-proton fluctuations $\kappa_4/\kappa_1$ across $\sqrt{s}$ \cite{ZHANGSQM} appearing flat.
Factorial cumulants show qualitative features  predicted from critical point models \cite{Bzdak:2019pkr}.
Viscous effects may smear out or alter critical fluctuations \cite{Dore:2022qyz} and the critical region can either cause critical lensing or minimize the effect \cite{Nonaka:2004pg,Dore:2022qyz,Pradeep:2024cca,Karthein:2024zvs}.
Theorists are developing frameworks (initial state, fluid dynamics, hadron interactions) that dynamically study $n_B\neq 0$ (and strangeness, $n_S\neq 0$, and electric charge, $n_Q\neq 0$). 
The Taylor series EOS from lattice is only valid up to  $\mu_B/T\sim [2-3]$ \cite{Bollweg:2022rps}, which does not cover the entire EOS needed for Beam Energy Scan II. 
An alternative expansion was developed that  improves the reach in $\mu_B/T$ \cite{Borsanyi:2021sxv} and is coupled to 3D Ising \cite{Kahangirwe:2024cny}.
In hydro simulations, every fluid cell must have a valid EOS.  In \cite{Plumberg:2024leb} a solution to a limited EOS was developed in 4-dimensions ($T,\mu_B,\mu_S,\mu_Q$) to smoothly map to a conformal EOS for out-of-bound fluid cells. 
The MUSES collaboration   provides open-source EOS across the phase diagram \cite{MUSESalpha}.

Initial conditions either include baryon stopping through strings \cite{Shen:2017bsr}, a 3-fluid model \cite{Cimerman:2023hjw}, color glass condensate \cite{Garcia-Montero:2023gex,Garcia-Montero:2024jev}, or a hadron transport \cite{Schafer:2021csj}. 
It is also possible to study BSQ (baron, strangeness, electric charge) fluctuations where net-$n_B=0$ through gluon splittings into $q\bar{q}$ pairs \cite{Carzon:2019qja}.
These initial states may be far-from-equilibrium and require a pre-equilibrium state incorporating BSQ charges \cite{Carzon:2023zfp,DoreSQM}. 
They are fed into relativistic viscous BSQ hydrodynamic codes, existing open-source codes  are \texttt{CCAKE} \cite{Plumberg:2024leb}, \texttt{MUSIC} \cite{Monnai:2024pvy}, and \texttt{vHLLE} \cite{Karpenko:2013wva}, each have different equations of motion, numerical algorithms, and EOS solvers. 
Stability constraints \cite{Almaalol:2022pjc} provide further constraints on BSQ hydro.
Phase transitions \cite{Kapusta:2024riv} or critical fluctuations \cite{Karthein:2024ipl,An:2022jgc} incorporation in relativistic  hydrodynamics is underway but not yet implemented in dynamical frameworks. 
Hadronization also requires criticality  \cite{Pradeep:2022eil}.
For low $\sqrt{s}$ the influence of the QGP  may  short lived such that it is negligible \cite{Inghirami:2022afu}.  An alternative  is to incorporate potentials within a hadron transport to simulation phase transitions \cite{Sorensen:2020ygf}.


{\bf NS} Scientists can now extract posteriors of the NS EOS ($T=0$, large $n_B$, at $\beta$-equilibrium). 
Different approaches are used: functional forms, physics based EOS (often multiple stitched together in different ranges of $n_B$), or a mixture. 
There is strong evidence that the speed of sound squared surpasses the conformal limit i.e. $c_s^2>1/3$ to support heavy NS \cite{Bedaque:2014sqa} and may near the causal limit if ultra heavy NS exist \cite{Tan:2020ics,Kojo:2020krb}.
A bump in $c_s^2$ could indicate a cross-over  into quarks \cite{McLerran:2018hbz} or a phase transition into hyperons \cite{Cruz-Camacho:2024odu}.
Functional forms should  capture features in $c_s^2$ that indicate phase transitions/new degrees of freedom, which is not possible with polytropes/spectral EOS \cite{Tan:2020ics,Tan:2021ahl}. 
Non-parametric approaches can reproduce these features like Gaussian processes \cite{Essick:2023fso}, these features can also be built in  \cite{Mroczek:2023zxo}.
Strong phase transitions  could be measured via mass twins \cite{Alford:2013aca} or the binary love relation \cite{Tan:2021nat}. 
We cannot calculate the EOS directly at large $n_B$. 
However, lattice QCD calculations are possible at $T=0$ with isospin asymmetry, leading to $c_s^2>1/3$ \cite{Brandt:2022hwy}. 
pQCD calculations are relevant at $n_B\gtrsim 40\,n_{sat}$. 
Using causality and stability, one can obtain bounds that reach to smaller $n_B$ to further constrain the EOS \cite{Komoltsev:2021jzg}.
Bulk viscosity within NS mergers arises from weak interactions, affecting different stages of the merger \cite{Most:2021zvc}, and is influenced by the degrees of freedom \cite{Alford:2020pld} and $Y_Q$ \cite{Yang:2023ogo}. 
In \cite{Ripley:2023lsq} the first constraints on bulk viscous in the inspiral of NS mergers  was extracted, making tests possible with further data.

Groups use low-energy HIC flow data  to extract the low $T\geq 0$ EOS for $Y_Q\sim0.5$.
It is important to consider structure in $c_s^2$, as was done first in \cite{Oliinychenko:2022uvy}, indicating a large bump in $c_s^2$ around $[2,3]n_{sat}$. 
It is possible to convert a given NS EOS into HIC using the symmetry energy expansion (and applying saturation and causality/stability constraints) \cite{Yao:2023yda}, which indicated  ultraheavy NS are consistent with HIC.
Next is to expand cold EOS into finite $T$, which is possible using a $T/\mu_B$ expansion  \cite{Mroczek:2024sfp}  or an effective mass (npe matter) \cite{Raithel:2019gws}.

\section{Bulk Dynamics}
\label{sec:Hadrons}
The QGP is well described by a relativistic viscous fluid with small viscosity \cite{Heinz:2013th}, leading to significant developments of relativistic \emph{viscous} fluids. At $n_B=0$ and large systems, the properties can be exacted from a Bayesian analysis \cite{Bernhard:2019bmu,JETSCAPE:2020mzn,Nijs:2020roc}. 
However,  the QGP medium may begin extremely far-from-equilibrium such that even causality and stability break down \cite{Plumberg:2021bme}, which affects the extraction of bulk viscosity \cite{Domingues:2024pom}. 
Collisions of light nuclei pushes the boundaries of fluid dynamics even further \cite{PHENIX:2018lia}, spin hydrodynamics is relevant for $\Lambda$ polarization results \cite{STAR:2017ckg,Florkowski:2017ruc}, and magnetohydrodynamics  is needed to explore  large magnetic fields \cite{Gursoy:2014aka}.
\emph{Key Questions: } Does HIC knowledge of rel. viscous
fluids affect other fields? Could hydrodynamics be
relevant for the EIC? What role does spin and magnetic fields play?

{\bf Advances in fluid dynamics} Predictions between $\sqrt{s}$ \cite{Niemi:2015voa,Noronha-Hostler:2015uye} were confirmed \cite{ALICE:2016ccg}, so one could  extract nuclear structure information from collisions of deformed ions \cite{Bally:2022vgo}. 
One can extract  the neutron skin from $^{208}Pb$ \cite{Giacalone:2023cet} from HIC, which was more consistent with ab initio \cite{Hu:2021trw} than the PREXII data \cite{PREX:2021umo}.
Collaborations with nuclear structure physicists have led to suggestions of new  run to shed light on $\alpha$ clustering \cite{Summerfield:2021oex} or interesting deformations \cite{Giacalone:2024luz,Giacalone:2024ixe}.
Fundamental questions still remain for small systems and peripheral collisions. Correlations between $v_2$ and $\langle p_T\rangle$ are hard to reproduce \cite{ALICE491854}.
In extremely small systems i.e. a vector meson-ion collisions (ultra peripheral collisions) there are hints of collective behavior \cite{Zhao:2022ayk}, which has interesting implications for the Electron Ion Collider. 
These advanced have led to cold atoms connections \cite{Floerchinger:2021ygn}.

{\bf Spin and Magnetic Fields}
HIC have extremely large, short lived magnetic fields \cite{Inghirami:2019mkc}. Groups are working on developing relativistic magnetohydrodyanmics (MHD) codes for HIC. 
Since ideal MHD is standard in astrophysics, at least two astro codes have been converted to HIC \cite{Inghirami:2016iru,Mayer:2024kkv}.   HIC physicists  and astrophysicists incorporated Israel-Stewart  equations of motion into astro MHD codes \cite{Most:2021rhr}.
Non-linear stability and causality techniques from HIC \cite{Bemfica:2020xym}, have been extended to MHD coupled to general relativity for accretion disks around black holes \cite{Cordeiro:2023ljz}.
Similar constraints are also available for spin/chiral rel. hydrodynamics \cite{Abboud:2023hos,Daher:2024bah}.
$\Lambda$ polarization can further constrain hydrodynamic parameters \cite{PalermoSQM}.

\section{Hadron Physics}
\label{sec:Hadrons}
HIC cannot directly probe quarks and gluons, rather detectors measure  charged hadrons. 
Understanding their properties provides insight into  freeze-out, hadronization, and  Brownian motion.
\emph{Key Questions:} How is strangeness produced?  What causes isospin breaking? Can we understand heavy flavor across system size? 
Could charm quarks be thermalized?

{\bf Puzzles from hadrons} The hadron resonance gas (HRG) model helped understand hadronization.  
HRG is used to extract freeze-out using thermal fits \cite{Andronic:2017pug} and  fluctuations \cite{Bellwied:2018tkc,Alba:2020jir} (see also in lattice QCD \cite{Borsanyi:2014ewa}). 
A tension remains between light and strange hadrons (see \cite{Floris:2014pta}).
Possibly increasing the number of resonances would solve the puzzle \cite{Bazavov:2014xya}, but the tension remained even with new states \cite{Alba:2020jir,SanMartin:2023zhv}.
Either there are separate freeze-out temperatures  \cite{Bellwied:2013cta} or an S-matrix approach is required \cite{Andronic:2018qqt}. 
Extensions to the HRG include magnetic fields \cite{Vovchenko:2024wbg} or surface tension \cite{ZherebtsovaSQM}.
Experiments found isospin symmetry breaking of kaons across $\sqrt{s}$ \cite{Brylinski:2023nrb}. 
One explanation  is  a disorientated isospin condensate \cite{Kapusta:2023xrw}.

{\bf Puzzles from heavy flavor} Heavy flavor  provides an interesting probe of the QGP that is normally modeled through a Langevin equation at low  $p_T$, or an energy loss model at high $p_T$.
Most dynamics are described using more weakly coupled approaches \cite{BassSQM}, but others have used charm as a conserved charge within the fluid \cite{Capellino:2022nvf,Capellino:2023cxe}.
Coalescence is important at low $p_T$  to correctly capture the hadronization process (see \cite{HirayamaSQM} for hadronic rescattering). 
While it was thought heavy flavor models lead to disparate results for the nuclear modification factors $R_{AA}$,  much could be attributed to different medium effects since the same medium led to very similar results \cite{Cao:2018ews}. 
After the first event-by-event studies with heavy flavor \cite{Nahrgang:2014vza} and jets \cite{Noronha-Hostler:2016eow}, soft-heavy \cite{Prado:2016szr,Sambataro:2022sns} and soft-hard \cite{Barreto:2022ulg,Holtermann:2023vwr} correlations advanced. 
These correlations can constrain the $T$ dependence of diffusion \cite{Sambataro:2023tlv}.
ALICE experimental data showed that the heavy flavor elliptical flow, $v_2$, scales with the soft $v_2$ \cite{Sambataro:2022sns}. 
Charmonium flow may be more complicated due to its internal structure \cite{Cho:2023kpe}.
Jets' sensitivity to event-by-event fluctuations can be seen by experimental measurements of jet $v_3\neq =0$ \cite{ATLAS:2021ktw}.
Groups are trying to unify the soft-heavy \cite{Zhao:2024ecc} and soft-hard \cite{Karmakar:2024jak} frameworks.

Small systems have  perplexing behavior for  heavy flavor. $R_{AA}\sim 1$, while  pPb D meson $v_2$ is large  
\cite{CMS:2018loe} but charmonium \cite{CMS:2018duw} and bottomonium $v_2\sim 0$ \cite{CMS:2023dse} and in pp D mesons  $v_2>0$  and B mesons $v_2<0$ \cite{ATLAS:2023vms}.
A heavy flavor system size scan was proposed \cite{Katz:2019fkc}  with  multicharm hadrons predictions in \cite{Minissale:2023dct} . 
Corrections to the short path length may be required \cite{Faraday:2023mmx}.
Heavy flavor requires longer formation times \cite{Moore:2004tg} due to their masses. 
The interplay between the early far-from-equilibrium stage and heavy flavor is intriguing, such as how $c\bar{c}$ and $b\bar{b}$ pairs disassociate in the glasma \cite{Pooja:2024rnn}.

{\bf Acknowledgments} J.N.H thanks  SQM24 for providing child care and acknowledges support from DOE Grant No. DE-SC0023861,  Saturated Glue (SURGE) Topical Theory Collaboration, and NSF Grant No. OAC-2103680  within the framework of MUSES collaboration.

\bibliography{inspire,NOTinspire}

\end{document}